if the sides of a square are lengthned by 7 inches,the area becomes 121 in^2. find the length of a side of the original square
Answer by oberobic(2304)

s = lengths of the sides of the square
s^2 = area
(s+7)^2 = 121 given
(s+7)(s+7) = 121
s^2 +7s +7s + 49 = 121
.
s^2 +7s +7s + 49 -121 = 0
s^2 +14s -72 = 0
.
factor
(s + ?)(s - ?) = 0
.
2*36 = 72, difference = 34, nope.
4*18 = 72, difference = 14, which is it!
.
(s +18)(s -4) = 0
.
s = -18 or 4
a negative length does not make sense, so s = 4
.
add 7 and check area to make sure it is 121
.
11*11=121 is true.
.
The sides of the original square were 4 inches.
.
Done.